# Human-in-the-Loop Feature Selection Using Interpretable Kolmogorov-Arnold Network-based Double Deep Q-Network

### v1

@@ Abstract
Feature selection is critical for improving the performance and interpretability of machine learning models, particularly in high-dimensional spaces where complex feature interactions can reduce accuracy and increase computational demands. Existing approaches often rely on static feature subsets or manual intervention, limiting adaptability and scalability. However, dynamic, per-instance feature selection methods and model-specific interpretability in reinforcement learning remain underexplored. This study proposes a human-in-the-loop (HITL) feature selection framework integrated into a Double Deep Q-Network (DDQN) using a Kolmogorov-Arnold Network (KAN). Our novel approach leverages simulated human feedback and stochastic distribution-based sampling, specifically Beta, to iteratively refine feature subsets per data instance, improving flexibility in feature selection. The KAN-DDQN achieved notable test accuracies of 93% on MNIST and 83% on FashionMNIST, outperforming conventional MLP-DDQN models by up to 9%. The KAN-based model provided high interpretability via symbolic representation while using 4 times fewer neurons in the hidden layer than MLPs did. Comparatively, the models without feature selection achieved test accuracies of only 58% on MNIST and 64% on FashionMNIST, highlighting significant gains with our framework. Pruning and visualization further enhanced model transparency by elucidating decision pathways. These findings present a scalable, interpretable solution for feature selection that is suitable for applications requiring real-time, adaptive decision-making with minimal human oversight.
Index Terms:
@@
Human-in-the-Loop, Feature Selection, Kolmogorov Arnold Network, Reinforcement Learning, Double Deep Q-Network.
I Introduction
@@
Feature selection is essential for building efficient, accurate, and robust machine learning models [1, 2]. While models ideally should automatically identify the most predictive features, a high-dimensional input space can significantly hinder performance, often requiring large volumes of data to learn from the complex relationships between features effectively. This phenomenon, known as the “curse of dimensionality,” increases computation time and resource use. Consequently, selecting a smaller subset of relevant features improves performance and makes the model more cost-effective.
@@
One common solution is to incorporate expert knowledge to determine the most useful features; however, this process is costly, time-consuming, and highly manual. Additionally, experts with deep domain knowledge are often not involved in the actual design and development of the model. Automatic feature selection methods offer an alternative, ranking features by their relevance or importance; however, both manual and automatic approaches typically yield a single subset of features for the entire dataset, which may not capture variability across individual observations. When training data are sparse relative to the feature space, the use of a single subset can limit the model’s ability to generalize effectively across all instances [3].
@@
To address the challenge of per-example feature selection, we propose a reinforcement learning (RL) framework that leverages simulated feedback to replicate human feature selection. Our approach employs a Double Deep Q-Network (DDQN) setup with a Kolmogorov-Arnold Network (KAN) as both the Q and target networks. This model-specific interpretable KAN-based RL structure aims to refine feature selection on a per-example basis iteratively. In this setup, simulated feedback, rather than direct human input, is a proxy for expert annotation. The feedback signal highlights the most relevant features for each data example, creating a pathway for the model to prioritize these features during training. Unlike prior methods, our model explores distributions beyond Bernoulli and organizes convolutional and pooling layers to ensure feature maps match the input and simulated feedback shapes. RL then optimizes a policy to select a unique subset of features per observation. By minimizing the classifier’s prediction loss and the discrepancy between simulated feedback and the model-selected features, this policy yields feature subsets that improve the interpretability and performance of the final predictions. Since predictions are based only on the selected feature subsets, this method offers interpretable, case-specific insights into the model’s output. Using simulated feedback further enables the model to reflect causal structures likely to be relevant in practical applications. We validate our methodology through rigorous experimentation on benchmark datasets, showing the efficiency of our approach in improving model accuracy while maintaining computational feasibility. We aim to establish best practices for integrating human feedback into the feature selection process by investigating the influence of various hyperparameters, stochastic distributions, and the absence of feature selection.
@@
    Our research presents a simulated feedback mechanism that generates feature relevance feedback without human annotators, facilitating a scalable training process that reflects the causal relationships typically identified by human experts.
@@
The following sections provide an overview of our work: Section II reviews the relevant background; Section III outlines the complete methodology of our proposed framework; Section IV details the experimental design; Section V discusses the results and their interpretations; and finally, Section VI concludes the research while outlining potential future directions.
II Background
II-A Human-in-the-Loop Feature Selection Using RL
@@
Feature selection is critical in developing machine learning models but is often executed through data-driven methods that overlook insights from human designers [2, 4]. We introduce a HITL [5] framework that integrates simulated feedback to identify the most relevant variables for specific tasks, which can be modeled using DDQN-based [6] RL to facilitate per-example feature selection [7], enabling the model to minimize its loss function while emphasizing significant variables from a simulated human perspective. A major gap in RL is the limited model-specific interpretability, as most models operate as black-box systems, making it challenging to understand their decision-making processes [8]. Our KAN agent enhances the interpretability of DDQN-based RL by providing symbolic representations of learned policies. Our methodology employs variable elimination techniques, focusing on selecting subsets of features rather than using embedded methods. This approach optimizes feature selection and learning processes concurrently via gradient descent and distinguishes itself by selecting different subsets for each observation. This enhances interpretability, as the chosen variables represent the “causes” driving the model’s predictions. While previous work [1, 5] has explored per-example feature selection via traditional filter methods based on mutual information, our framework goes beyond incorporating simulated human feedback for a more dynamic approach. Traditional feature selection techniques include filter methods, which select top-ranked features based on criteria like mutual information; wrapper methods, which evaluate subsets of features by retraining models for each subset; and embedded methods, which attempt to select features during the model training process. Unlike these approaches, which can be sequential and computationally costly, our model generates candidate subsets in a single step, guided by simulated feedback, thereby avoiding arbitrary stopping criteria and allowing for real-time, per-example selection in complex datasets. Our work draws inspiration from the probabilistic knowledge elicitation [9, 10] focused on querying users for global feature relevance. However, our model goes beyond the focus on Bernoulli distributions [3] by exploring various distributions, including beta, Gaussian, and Gumbel-Softmax, enhancing the flexibility and effectiveness of feature selection. This breadth allows us to capture a wider range of relationships within the data compared with previous approaches.
II-B Kolmogorov-Arnold Networks (KANs)
@@
The differences between KAN and MLP are that KANs implement learnable activation functions along edges, contrasting with MLPs, which implement fixed activation functions at nodes. KANs represent weights as splines, improving their ability to approximate complex functions with fewer parameters. KANs and MLPs can be extended to multiple layers, supporting deep architectures.
III Methodology
Input: Input shape $input\_shape$, Number of filters $num\_filters$, Number of convolutional layers $num\_conv\_layers$, Hidden layer dimension $hidden\_dim$, Distribution $distribution$, Temperature parameter $\tau$
@@ III Methodology
2 Function FSNet($input\_shape$, $num\_filters$, $num\_conv\_layers$, $hidden\_dim$, $distribution$, $\tau$):
3       Initialize convolutional layer parameters: $channels\leftarrow$ []; for $i\leftarrow 1$ to $num\_conv\_layers$ do
@@ 4             $num\_filters\leftarrow num\_filters\times 2$;
5             Append $num\_filters$ to $channels$;
6
7      Build feature extractor with $num\_conv\_layers$ Conv2D, ReLU, and MaxPool layers;
8       Initialize fully connected layer $fc1$ with input size $n\_features$ and hidden size $hidden\_dim$;
9       Initialize output layer $fc2$ with input size $hidden\_dim$ and output size $n\_features$;
10       Initialize activation functions: $relu$, $sigmoid$, and $softmax$;
11
12
13 Function forward($x$, $feedback$, $epoch$):
14       Reshape $x$ to match $input\_shape$;
@@ 18       Compute feature selection probabilities $probs\leftarrow sigmoid(logits)$;
19
20      if $feedback$ is None then
@@ 21             $feedback\leftarrow probs$;
22
23
24      Sample $sample\_probs$ from $distribution$ using $logits$ as parameters;
@@ 28       $selected\_feature\leftarrow features\times sample\_probs$;
29
30      return $selected\_feature$, $feedback\_cost$, $probs$;
31
@@
III-A Human-in-the-Loop Feature Selection Process
The feature selection network (FSNet) (see Algorithm 1) is designed to determine the relevance of each input feature for a given task. The architecture combines convolutional layers for feature extraction with various probabilistic distributions to model and sample feature importance. Here, we detail each model component, focusing on architectural elements, distribution-based sampling, and feedback alignment.
III-A1 Convolutional Feature Extraction Layers
@@
where $F^{(i)}\in\mathbb{R}^{f_{i}\times\frac{h}{2^{i}}\times\frac{w}{2^{i}}}$ and $F^{(0)}=x$. Each convolutional layer uses a $3\times 3$ kernel and padding of 1 to maintain spatial dimensions, while max-pooling layers with kernel size 2 reduce the height and width by half at each step.
III-A2 Fully Connected Layers for Feature Probability Mapping
@@
III-A3 Distribution-Based Sampling for Probabilistic Feature Selection
In our framework, feature selection is modeled as a probabilistic policy $\pi(a\mid\hat{q})$, where each feature selection mask $a\in\{0,1\}^{d}$ is sampled based on a probability vector $\hat{q}=(\hat{q}_{1},\hat{q}_{2},\ldots,\hat{q}_{d})$. This vector $\hat{q}$, derived from the network output, represents the relevance probability of each feature. For a given action $a$ (or feature mask), the policy is defined as:
@@ III-A3 Distribution-Based Sampling for Probabilistic Feature Selection
where $P(a_{j}\mid\hat{q}_{j})$ is the probability of selecting (or discarding) each feature $j$ based on a chosen distribution. Using this probabilistic policy allows flexibility in feature selection while controlling each selection probability by the network’s learned relevance score $\hat{q}_{j}$. The feature mask $a$ is generated according to various probabilistic distributions described below.
Bernoulli Distribution Sampling: In Bernoulli distribution sampling, each feature’s selection is modeled as an independent binary random variable, where each feature $j$ is selected with a probability $\hat{q}_{j}$ derived from the network’s sigmoid output.
$a_{j}\sim\text{Bernoulli}(\hat{q}_{j})$ (8)
where $\hat{q}_{j}\in[0,1]$ represents the probability of selecting feature $j$. If $a_{j}=1$, feature $j$ is retained; if $a_{j}=0$, feature $j$ is discarded.
Gumbel-Softmax Sampling: Gumbel-Softmax distribution provides a differentiable approximation to categorical sampling. For each feature $j$, the mask $a_{j}$ is computed as:
$a_{j}=\frac{\exp((\log\hat{q}_{j}+g_{j})/\tau)}{\sum_{k=1}^{d}\exp((\log\hat{q}_{k}+g_{k})/\tau)}$ (9)
where $g_{j}\sim\text{Gumbel}(0,1)$ and the temperature $\tau$ decays over time.
Gaussian Distribution Sampling: Each feature selection probability $\hat{q}_{j}$ is used as the mean $\mu_{j}$ of a Gaussian distribution with fixed standard deviation $\sigma=1$:
$a_{j}\sim\mathcal{N}(\mu_{j}=\hat{q}_{j},\sigma=1)$ (10)
This sampling scheme maintains smooth variations around $\hat{q}_{j}$.
Beta Distribution Sampling: Beta distribution, bounded between 0 and 1, provides flexibility in modeling the probability $\hat{q}_{j}$ with shape parameters $\alpha_{j}$ and $\beta_{j}$:
$a_{j}\sim\text{Beta}(\alpha_{j},\beta_{j}),\quad\alpha_{j}=\beta_{j}=\text{softplus}(\hat{q}_{j})+1$ (11)
Beta distribution allows for high flexibility in the shape of the probability distribution, which can vary significantly depending on $\hat{q}_{j}$. For example, if $\hat{q}_{j}$ is close to 1, the distribution skews towards retaining the feature; if $\hat{q}_{j}$ is close to 0, it skews towards discarding it.
Dirichlet Distribution Sampling: For interdependent feature probabilities, we generalize to the Dirichlet distribution, where each $\hat{q}_{j}$ represents a concentration parameter:
$\mathbf{a}\sim\text{Dirichlet}(\alpha),\quad\alpha_{j}=\text{softplus}(\hat{q}_{j})+1$ (12)
where $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$ is the vector of concentration parameters.
Multinomial Distribution Sampling: In multinomial sampling, we interpret $\hat{q}=(\hat{q}_{1},\hat{q}_{2},\ldots,\hat{q}_{d})$ as logits:
$\mathbf{a}\sim\text{Multinomial}(n=1,p=\sigma(\hat{q}))$ (13)
where $p$ is the probability distribution over the features given by applying softmax to $q$ and $n=1$ indicates that we are drawing a single sample. This approach allows for categorical selection from among the features based on the probability distribution defined by $q$.
Laplace Distribution Sampling: In Laplace sampling, each $\hat{q}_{j}$ serves as the mean $\mu_{j}$ for the Laplace distribution:
$a_{j}\sim\text{Laplace}(\mu_{j}=\hat{q}_{j},\text{scale}=1)$ (14)
Cauchy Distribution Sampling: For Cauchy distribution sampling, each $\hat{q}_{j}$ is the location parameter:
$a_{j}\sim\text{Cauchy}(\text{location}=\hat{q}_{j},\text{scale}=1)$ (15)
The Cauchy distribution has no well-defined mean or variance, making it highly sensitive to extreme values. This can result in selective or random feature activation depending on the value of $\hat{q}_{j}$, which introduces higher variability in feature selection.
Uniform Distribution Sampling: Uniform sampling considers each $a_{j}$ equally probable within $[0,1]$:
$a_{j}\sim\text{Uniform}(0,1)$ (16)
The sampled feature mask $\mathbf{a}$ is then applied to the feature vector $\mathbf{f}$ as:
$\mathbf{f}_{\text{selected}}=\mathbf{f}\odot\mathbf{a}$ (17)
where $\odot$ denotes element-wise multiplication, yielding a filtered feature vector $\mathbf{f}_{\text{selected}}$ that selectively retains features based on their relevance.
III-A4 Feedback Alignment with MSE Loss

### v2

@@ Abstract
Feature selection is critical for improving the performance and interpretability of machine learning models, particularly in high-dimensional spaces where complex feature interactions can reduce accuracy and increase computational demands. Existing approaches often rely on static feature subsets or manual intervention, limiting adaptability and scalability. However, dynamic, per-instance feature selection methods and model-specific interpretability in reinforcement learning remain underexplored. This study proposes a human-in-the-loop (HITL) feature selection framework integrated into a Double Deep Q-Network (DDQN) using a Kolmogorov-Arnold Network (KAN). Our novel approach leverages simulated human feedback and stochastic distribution-based sampling, specifically Beta, to iteratively refine feature subsets per data instance, improving flexibility in feature selection. The KAN-DDQN achieved notable test accuracies of 93% on MNIST and 83% on FashionMNIST, outperforming conventional MLP-DDQN models by up to 9%. The KAN-based model provided high interpretability via symbolic representation while using 4 times fewer neurons in the hidden layer than MLPs did. Comparatively, the models without feature selection achieved test accuracies of only 58% on MNIST and 64% on FashionMNIST, highlighting significant gains with our framework. We further validate scalability on CIFAR-10 and CIFAR-100, achieving up to 30% relative macro F1 improvement on MNIST and 5% on CIFAR-10, while reducing calibration error by 25%. Complexity analysis confirms real-time feasibility with latency below 1 ms and parameter counts under 0.02M. Pruning and visualization further enhanced model transparency by elucidating decision pathways. These findings present a scalable, interpretable solution for feature selection that is suitable for applications requiring real-time, adaptive decision-making with minimal human oversight. The code is available at https://github.com/Abrar2652/HITL-FS.
{IEEEkeywords}
@@
Human-in-the-Loop, Feature Selection, Kolmogorov-Arnold Network, Reinforcement Learning, Double Deep Q-Network.
1 Introduction
[…]
@@ 1 Introduction
Feature selection is essential for building efficient, accurate, and robust machine learning models [6, 2]. While models ideally should automatically identify the most predictive features, a high-dimensional input space can significantly hinder performance, often requiring large volumes of data to effectively learn the complex relationships between features. This phenomenon, known as the “curse of dimensionality,” increases computation time and resource use. Consequently, selecting a smaller subset of relevant features improves performance and makes the model more cost-effective.
@@
One common solution is to incorporate expert knowledge to determine the most useful features; however, this process is costly, time-consuming, and highly manual. Additionally, experts with deep domain knowledge are often not involved in the actual design and development of the model. Automatic feature selection methods offer an alternative, ranking features by their relevance or importance; however, both manual and automatic approaches typically yield a single subset of features for the entire dataset, which may not capture variability across individual observations. When training data are sparse relative to the feature space, using a single subset can limit the model’s ability to generalize effectively across all instances [3].
@@
To address the challenge of per-example feature selection, we propose a reinforcement learning (RL) framework that leverages simulated feedback to replicate human feature selection. Our approach employs a Double Deep Q-Network (DDQN) setup with a Kolmogorov-Arnold Network (KAN) [8, 5] as both the Q-network and the target network. This model-specific, interpretable KAN-based RL structure aims to refine feature selection on a per-example basis iteratively. In this setup, simulated feedback, rather than direct human input, is a proxy for expert annotation [1, 4]. The feedback signal highlights the most relevant features for each data example, enabling the model to prioritize these features during training. Unlike prior methods [3], our model explores distributions beyond Bernoulli and organizes convolutional and pooling layers to ensure feature maps match the input and simulated feedback shapes. RL then optimizes a policy to select a unique subset of features per observation. By minimizing the classifier’s prediction loss and the discrepancy between simulated feedback and the model-selected features, this policy yields feature subsets that improve the interpretability and performance of the final predictions. Since predictions are based only on the selected feature subsets, this method offers interpretable, case-specific insights into the model’s output. Using simulated feedback further enables the model to reflect causal structures likely to be relevant in practical applications. We validate our methodology through rigorous experimentation on benchmark datasets, showing the efficiency of our approach in improving model accuracy while maintaining computational feasibility. We aim to establish best practices for integrating human feedback into the feature selection process by investigating the influence of various hyperparameters, stochastic distributions, and the absence of feature selection.
In many deployed classification settings, such as safety triage in vision, early exits in mobile perception, and clinician-in-the-loop screening, models must meet tight memory/latency budgets while preserving case-wise interpretability. Static, dataset-level feature subsets often underperform when intra-class variability is high; conversely, per-instance selection can reduce redundant computation and expose the causal drivers of individual decisions. Our goal is to operationalize instance-wise feature selection under realistic constraints: (i) a compact head suitable for resource-limited environments, (ii) a stochastic, differentiable gating mechanism that trades sparsity for accuracy, and (iii) a reinforcement-learning controller whose policy remains interpretable. This motivates our HITL-DDQN framework, which features a KAN or MLP head and a stochastic gate aligned with simulated feedback, thereby delivering actionable sparsity–accuracy operating points and model-specific interpretability.
@@
    Our research presents a simulated feedback mechanism that generates feature relevance feedback without the need for human annotators, facilitating a scalable training process that reflects the causal relationships typically identified by human experts.
@@
The following sections provide an overview of our work: Section 2 reviews the relevant background; Section 3 outlines the complete methodology of our proposed framework; Section 4 details the experimental design; Section 5 discusses the results and their interpretations; and finally, Section 6 concludes the research while outlining potential future directions.
2 Background
2.1 Human-in-the-Loop Feature Selection Using RL
@@
Feature selection is critical in developing machine learning models but is often executed through data-driven methods that overlook insights from human designers [2, 9]. We introduce a HITL [16] framework that integrates simulated feedback to identify the most relevant variables for specific tasks, which can be modeled using DDQN-based [7] RL to facilitate per-example feature selection [13], enabling the model to minimize its loss function while emphasizing significant variables from a simulated human perspective. A major gap in RL is the limited model-specific interpretability, as most models operate as black-box systems, making it challenging to understand their decision-making processes [14]. Our KAN agent enhances the interpretability of DDQN-based RL by providing symbolic representations of learned policies. Our methodology employs variable elimination techniques, focusing on selecting subsets of features rather than using embedded methods. This approach optimizes feature selection and learning processes concurrently via gradient descent, distinguishing itself by selecting different subsets for each observation. This enhances interpretability, as the chosen variables represent the “causes” driving the model’s predictions. While previous work [6, 16] has explored per-example feature selection via traditional filter methods based on mutual information, our framework goes beyond incorporating simulated human feedback for a more dynamic approach. Traditional feature selection techniques include filter methods, which select top-ranked features based on criteria like mutual information; wrapper methods, which evaluate subsets of features by retraining models for each subset; and embedded methods, which attempt to select features during the model training process. Unlike these approaches, which can be sequential and computationally costly, our model generates candidate subsets in a single step, guided by simulated feedback, thereby avoiding arbitrary stopping criteria and allowing for real-time, per-example selection in complex datasets. Our work draws inspiration from the probabilistic knowledge elicitation [1, 15] focused on querying users for global feature relevance. However, our model goes beyond the focus on Bernoulli distributions [3] by exploring various distributions, including beta, Gaussian, and Gumbel-Softmax, enhancing the flexibility and effectiveness of feature selection. This breadth allows us to capture a wider range of relationships within the data compared with previous approaches.
2.2 Kolmogorov-Arnold Networks (KANs)
@@
The primary difference between KAN and MLP is that KANs implement learnable activation functions along edges, whereas MLPs implement fixed activation functions at nodes. KANs represent weights as splines, improving their ability to approximate complex functions with fewer parameters. KANs and MLPs can be extended to multiple layers, supporting deep architectures. We introduce KANs here to frame their conceptual differences from MLPs; the practical instantiation inside our DDQN head (width, grid, regularization) is given later in Sec. 3-3.3 to avoid duplication and to keep implementation choices distinct from background theory.
Figure 1: System overview. The feature-selection head (FSNet) maps an input image to per-feature probabilities via a differentiable stochastic gate aligned to simulated feedback. The selected features are fed to a DDQN head (KAN or MLP). Replay-based training updates the Q-network, with periodic target sync. The gate induces instance-wise sparsity and exposes feature-level rationale, while the KAN head provides model-specific interpretability.
3 Methodology
Input: Input shape input_shapeinput\_shape, Number of filters num_filtersnum\_filters, Number of convolutional layers num_conv_layersnum\_conv\_layers, Hidden layer dimension hidden_dimhidden\_dim, Distribution distributiondistribution, Temperature parameter $\tau$
@@ 3 Methodology
21ex
3 Function FSNet(input_shapeinput\_shape, num_filtersnum\_filters, num_conv_layersnum\_conv\_layers, hidden_dimhidden\_dim, distributiondistribution, $\tau$):
4  Initialize convolutional layer parameters: $channels\leftarrow$ []; for $i\leftarrow 1$ to num_conv_layersnum\_conv\_layers do
@@ 5     $num\_filters\leftarrow num\_filters\times 2$;
6     Append num_filtersnum\_filters to channelschannels;
7
8  end for
9 Build feature extractor with num_conv_layersnum\_conv\_layers Conv2D, ReLU, and MaxPool layers;
10  Initialize fully connected layer $fc1$ with input size n_featuresn\_features and hidden size hidden_dimhidden\_dim;
11  Initialize output layer $fc2$ with input size hidden_dimhidden\_dim and output size n_featuresn\_features;
12  Initialize activation functions: relurelu, sigmoidsigmoid, and softmaxsoftmax;
13
14
15
161ex
17 Function forward($x$, feedbackfeedback, epochepoch):
18  Reshape $x$ to match input_shapeinput\_shape;
@@ 22  Compute feature selection probabilities $probs\leftarrow sigmoid(logits)$;
23
24 if feedbackfeedback is None then
@@ 25     $feedback\leftarrow probs$;
26
27  end if
28
29 Sample sample_probssample\_probs from distributiondistribution using logitslogits as parameters;
@@ 33  $selected\_feature\leftarrow features\times sample\_probs$;
34
35 return selected_featureselected\_feature, feedback_costfeedback\_cost, probsprobs;
36
37 return
@@
3.1 Human-in-the-Loop Feature Selection Process
The feature selection network (FSNet) (see Algorithm 1 and Fig. 1) is designed to determine the relevance of each input feature for a given task [13]. The architecture combines convolutional layers for feature extraction with various probabilistic distributions to model and sample feature importance. Here, we detail each model component, focusing on architectural elements, distribution-based sampling, and feedback alignment.
3.1.1 Convolutional Feature Extraction Layers
@@
where $F^{(i)}\in\mathbb{R}^{f_{i}\times\frac{h}{2^{i}}\times\frac{w}{2^{i}}}$ and $F^{(0)}=x$. Each convolutional layer uses a $3\times 3$ kernel and padding of 1 to maintain spatial dimensions, while max-pooling layers with a kernel size of 2 reduce the height and width by half at each step.
3.1.2 Fully Connected Layers for Feature Probability Mapping
@@
3.1.3 Distribution-Based Sampling for Probabilistic Feature Selection
[…]
@@ 3.1.3 Distribution-Based Sampling for Probabilistic Feature Selection
where $P(a_{j}\mid\hat{q}_{j})$ denotes the probability of selecting (or discarding) feature $j$ conditioned on its relevance score $\hat{q}_{j}$. This probabilistic formulation introduces flexibility by allowing the network to dynamically control the inclusion probability of each feature through its learned relevance estimate.
To parameterize the feature selection probabilities, we employ the Beta distribution, a continuous distribution bounded within $[0,1]$, which is well-suited for modeling probabilities, as shown in Table 1. Each feature’s selection probability is sampled as $a_{j}\sim\text{Beta}(\alpha_{j},\beta_{j})$, where the shape parameters are computed as:
$\alpha_{j}=\beta_{j}=\text{softplus}(\hat{q}_{j})+1$ (8)
Here, $\text{softplus}(x)=\ln(1+e^{x})$ ensures positivity and numerical stability for the shape parameters. The symmetry $\alpha_{j}=\beta_{j}$ allows the distribution to smoothly vary from uniform to highly skewed based on the learned relevance score $\hat{q}_{j}$. When $\hat{q}_{j}$ is large, the Beta distribution skews toward 1 (favoring feature retention), and when small, it skews toward 0 (favoring feature exclusion). This approach provides a highly flexible and differentiable mechanism for stochastic feature selection.
The sampled feature mask $\mathbf{a}$ is then applied to the feature vector $\mathbf{f}$ as $\mathbf{f}_{\text{selected}}=\mathbf{f}\odot\mathbf{a}$, where $\odot$ denotes element-wise multiplication, yielding a filtered feature vector $\mathbf{f}_{\text{selected}}$ that selectively retains features based on their relevance.
3.1.4 Feedback Alignment with MSE Loss